A company wants to give gifts to the top 1% of customers. If the mean spend per customer is $145 with a standard deviation of 35, what balance should the company specify?
Answer #1

Mean = 145, standard deviation = 35

For top 1% z = 2.326

x = mean + z*sd

x = 145 + 2.326*35 = 226.41

Hence, company should specify $226.41 amount as gift.